40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 13 19:02:56 2016, Teng, Update, General, ITMX sensor 47.png34.png
    Reply  Wed Sep 14 09:33:54 2016, Steve, Update, General, ITMX magnets are stucked ITMXmagStuck.png
       Reply  Wed Sep 14 20:05:32 2016, Lydia, Update, General, ITMX magnets are stucked 
Message ID: 12489     Entry time: Tue Sep 13 19:02:56 2016     Reply to this: 12491
Author: Teng 
Type: Update 
Category: General 
Subject: ITMX sensor 

[Lydia,Teng]

Something strange happened to the ITMX osem reading around 4.pm. PDT as shown below.

Also the there was no response of the reading as we adjusted the PITCH and YAW. :(

Note that we restarted the slow machine: c1susaux,c1ausex this afternoon because of the unresponced interface.

 

 

Attachment 1: 47.png  34 kB  | Hide | Hide all
47.png
Attachment 2: 34.png  18 kB  Uploaded Tue Sep 13 20:05:42 2016  | Hide | Hide all
34.png
ELOG V3.1.3-